# Thermoelectric effects of an interacting hadron gas in the presence of an external magnetic field

### v1

@@ Abstract
The hot and dense hadronic medium formed during the heavy-ion collisions at the RHIC and LHC energies can show thermoelectric effects in the presence of temperature gradients and non-zero baryon chemical potential. In this article we study the thermoelectric coefficients of an interacting hot and dense hadron gas using the relativistic Boltzmann transport equation (BTE) under the relaxation time approximation (RTA). We discuss the thermoelectric properties within different frameworks of Hardon Resonance Gas (HRG) models. In the presence of an external magnetic field, the thermoelectric coefficients become anisotropic, which leads to Hall-like thermoelectric coefficients, namely Nernst coefficients, along with the magneto-Seebeck coefficients. For the first time, we also estimate the Thomson coefficient of the medium, which comes into the picture due to the temperature dependence of the Seebeck coefficient of the medium. In the context of studying the thermoelectric generator performance, we calculate the values of the thermoelectric figure of merit of the medium.
@@ I Introduction
In heavy-ion collision experiments like the Large Hadron Collider (LHC) or the Relativistic Heavy Ion Collider (RHIC), a deconfined state of nuclear matter can be produced at extremely high temperatures ($T$) and/or baryon chemical potentials ($\mu_{B}$). This thermalized state of matter is called quark-gluon plasma (QGP). QGP evolves Brewer:2019oha ; Busza:2018rrf and cools down rapidly, producing a hot and dense hadronic matter that later freezes out leading the detection of long-lived final-state particles. Studying the hadronic phase followed by the hadronization process is crucial for understanding the full evolution of heavy-ion collisions and extracting hidden information about the properties of dense nuclear matter. Due to the extreme energy density in the initial phase, the system continues to expand and cool down rapidly. Hence, the temperature gradients develop both in space and time. The temperature decreases with an increase in spatial distance from the collision point, which leads to the radial temperature gradients. This temperature gradient plays a significant role in heavy-ion collisions because it, in turn, affects the evolution of the system. Currently, researchers are interested in understanding how the thermoelectric properties of the hot hadron gas influence the evolution of collision dynamics.
@@
The thermoelectric properties refer to the response of a material to a temperature gradient, which can generate an electric potential difference subjected to this temperature gradient book . A baryon current is produced in the medium at finite baryon chemical potential. Similarly, a charge current will also be produced due to the presence of charges. Due to the baryon and charge currents in the medium, thermal and electrical transport properties are of significant interest in the study of heavy-ion collisions Singh:2023pwf ; Singh:2023ues . The thermoelectric effects are proportional to the derivative of conductivity, and they are generally considered more sensitive to the responses to applied external fields. When thermoelectric coefficients are non-zero, temperature and chemical potential gradients in a conductor can drive an electric current. The Seebeck coefficient measures a material’s efficiency in converting these gradients into electric current and is defined under “open circuit” conditions, where the net current is zero Singh:2024emy . This coefficient has been widely studied in condensed matter systems, including superconductors, quantum dot-coupled Majorana bound states, and various junctions and materials PhysRevB.81.155457 ; PhysRevB.84.075420 ; PhysRevB.101.045101 ; PhysRevB.90.165115 . These thermoelectric transport properties of hot and dense hadron gas have been studied theoretically using various approaches such as relativistic kinetic theory Bhatt:2018ncr ; Das:2020beh ; Das:2021qii , hydrodynamics Greif:2017byw ; Prakash:1993bt ; Kadam:2014xka , and lattice QCD simulations Meyer_2011 ; Kaczmarek_2022 . For some situations, the Seebeck coefficient is not constant but shows a gradient with the temperature gradient. When a current drives through this gradient, the Thomson coefficient can be observed, which describes how much heat is absorbed or released. The amount of heat exchange depends on the medium’s properties and the direction of both the current and temperature gradient.
@@
In relativistic systems, the thermoelectric formalism differs notably from the nonrelativistic cases: here, thermal current requires a conserved number current, like the baryon number in QCD. This distinction affects the behavior of thermoelectric coefficients, as shown in studies of hadronic matter via the HRG model Das:2021qii . Studies in isotropic Abhishek:2020wjm and anisotropic Kurian:2021zyb ; Zhang:2020efz QGP systems have also explored this, although using perturbative QCD, which is more applicable at very high temperatures. Near the QCD transition temperature ($T_{c}$), nonperturbative effects are relevant, as the QCD vacuum structure is nontrivial, influenced by chiral symmetry breaking and confinement-related condensates. However, understanding the thermoelectric properties of the hadronic medium is also of equal importance. To understand the hot and dense hadronic matter formed in relativistic heavy-ion collisions, the ideal hadron resonance gas (IHRG) model is a widely used model. It explains the lattice QCD (lQCD) estimations of thermodynamic variables up to $T\approx 150~{}\mathrm{MeV}$ Bellwied:2013cta ; Bellwied:2017ttj , after which the hadrons start to melt, and the disagreements between the IHRG and lQCD appear. Unlike the lQCD, which breaks down at very high $\mu_{B}$ Borsanyi:2013bia ; HotQCD:2014kol , the IHRG model can be used to study highly dense hadronic or nuclear matter. However, the ideal HRG model is unable to explain the higher-order fluctuations of conserved charges, which have been predicted from the lQCD Bazavov:2013dta ; Bazavov:2017dus ; Borsanyi:2018grb . Thus, later on, better alternatives to the simplistic IHRG model have been explored extensively. One such widely used model is the excluded volume hadron resonance gas (EVHRG) model Andronic:2012ut . In this model, the hadrons have some finite radius and thus exert an outward pressure, which is the result of the excluded volume of the hadrons. EVHRG model explains the thermodynamic properties and the conserved charged fluctuations of the hadronic matter very well, giving a good agreement with the lQCD data Andronic:2012ut ; Kadam:2015xsa ; Kadam:2017iaz ; Pal:2020ucy . There are also other models, such as the repulsive mean-field hadron resonance gas (RMFHRG) model Kapusta:1982qd ; Olive:1980dy , which uses a potential term to introduce the repulsive interaction between the hadrons. Recently, a new model has been explored called the van der Waals hadron resonance gas (VDWHRG) model Vovchenko:2017cbu ; Vovchenko:2016rkn , which takes into account both attractive and repulsive interactions between the hadrons. It has been found that the VDWHRG model is the most reliable HRG model, which agrees with the lQCD estimations Vovchenko:2016rkn ; Samanta:2017yhh . It also gives a second-order liquid-gas phase transition critical point at very high $\mu_{B}$ Vovchenko:2015vxa , which can help study dense nuclear matter present in neutron stars.
@@
In this article, we offer a detailed examination of the thermoelectric properties and associated transport coefficients of the hot and dense hadronic medium both in the presence and absence of an external magnetic field. We have used four different formalisms, namely, the IHRG, EVHRG, RMFHRG and VDWHRG, to estimate the thermoelectric coefficients. Further, we also discuss the thermoelectric generator performance of the medium in terms of the thermoelectric figure of merit. This article is organized in the following manner. Sec. (II) briefly gives the derivation of thermoelectric coefficients with and without an external magnetic field. Sec. (III) discusses the results in detail and finally, in Sec. (IV), we summarize the study with possible outlooks.
@@ II Formalism
In this section, we discuss the hadron resonance gas model and its extended versions that we use to calculate the thermoelectric coefficients of hot and dense hadron gas. We calculate these coefficients both in the absence and presence of an external magnetic field by using the Boltzmann transport equation under relaxation time approximation. For all our calculations, we have taken all the hadrons and their resonances up to the mass cut-off of 2.6 GeV from the particle data group ParticleDataGroup:2008zun .
@@ II.1 Hadron Resonance Gas model
II.1.1 Ideal Hadron Resonance Gas Model (IHRG)
The IHRG model assumes that a system of hadrons and resonances behaves like an ideal gas, meaning that the particles do not interact with each other and are considered to be point particles. The grand canonical partition function for an ideal hadron resonance gas can be written as Pradhan:2022gbm ,
@@ II.1.1 Ideal Hadron Resonance Gas Model (IHRG)
Here, $g_{i}$ is the degeneracy, $k_{i}$ is momentum, $m_{i}$ is the mass and $\omega_{i}=\sqrt{k_{i}^{2}+m_{i}^{2}}$ is energy of the $i^{\rm th}$ hadron species. The $\pm$ sign corresponds to fermions and bosons, respectively. $\mu_{i}$ denotes the corresponding chemical potential, which is given by
@@
where baryon chemical potential is given by $\mu_{B}$ and $b_{i}$ denotes the baryon number of the $i^{\rm th}$ hadron. Here, we have considered a simple case by taking vanishing charge and strangeness chemical potential, i.e., $\mu_{Q}$ = $\mu_{S}$ = 0. The pressure $P_{i}$, energy density $\varepsilon_{i}$, and number density $n_{i}$ can now be obtained from the partition function, given as,
@@
II.1.2 Excluded Volume Hadron Resonance Gas Model (EVHRG)
The Excluded Volume Hadron Resonance Gas (EVHRG) model is an extension of the ideal Hadron Resonance Gas (IHRG) model that takes into account the finite size of hadrons. Meanwhile, in an IHRG model, hadrons are treated as point-like particles without considering their physical volume. The EVHRG model incorporates repulsive interactions by introducing an excluded volume for hadrons, preventing overlap and thereby providing a more realistic description of hot and dense hadronic matter created in heavy-ion collisions. In thermodynamically consistent excluded volume formulation, the transcendental equation for the pressure can be written as Rischke:1991ke ; Bhattacharyya:2013oya
@@ II.1.2 Excluded Volume Hadron Resonance Gas Model (EVHRG)
II.1.3 van der Waals Hadron Resonance Gas Model (VDWHRG)
@@
The total pressure in the VDWHRG model comes from the mesons, baryons, and anti-baryons separately because of the interaction and is given by Samanta:2017yhh ; Vovchenko:2016rkn
@@
Here, $M$, $B$, and $\bar{B}$ represent mesons, baryons, and anti-baryons. $\mu^{*M}$ is the modified chemical potential of mesons because of the excluded volume correction ($b_{M}$), and $\mu^{*B}$ and $\mu^{*\bar{B}}$ are the modified chemical potentials of baryons and anti-baryons due to VDW interactions (a and b) Sarkar:2018mbk . The modified chemical potential for mesons and (anti)baryons can be obtained from Eq. (2) and Eq. (15) as,
@@
where $n_{M}$, $n_{B}$ and $n_{\bar{B}}$ are the modified number densities of mesons, baryons and anti-baryons, respectively, which are given by
@@
II.1.4 Repulsive Mean-Field Hadron Resonance Gas Model (RMFHRG)
The RMF corrections are often implemented within a mean-field approximation, where the repulsive interactions are treated as an effective mean field acting on each hadron. This mean field modifies the single-particle energies of the hadrons, affecting their thermodynamic properties such as the equation of state, pressure, and density. A repulsive mean-field approach is utilized to incorporate the effects of repulsive interactions among hadrons, as it was used in Refs. Kapusta:1982qd ; Olive:1980dy and recently, in the case of baryons in Ref.Huovinen:2017ogf . In this approach, it is assumed that the repulsive interactions lead to a shift in the single particle energy and is given by
$\tilde{\omega}_{i}=\sqrt{k_{i}^{2}+m_{i}^{2}}+U(n)=\omega_{i}+U(n),$ (25)
@@ II.1.4 Repulsive Mean-Field Hadron Resonance Gas Model (RMFHRG)
In this work, we assign different repulsive interaction parameters for baryons and mesons. We denote the mean-field parameter for baryons ($B$) and anti-baryons ($\bar{B}$) by $K_{B}$, while for mesons we denote it by $K_{M}$. Thus, for baryons (antibaryons)
@@
where $n_{i}$ is the number density of $i^{\rm th}$ hadronic species. Note that $n_{B}$, $n_{\bar{B}}$, and $n_{M}$ are total baryon, antibaryon and meson number densities, respectively. Explicitly, for baryons,
@@
where the sum is over all the baryons. Here, $\mu_{\text{eff}}^{i}=b_{i}\mu_{B}-U(n_{B})$ is the baryon effective chemical potential, with $b_{i}$ being the baryonic number of $i^{\rm th}$ baryon and $\mu_{B}$, the baryon chemical potential. Similarly, for antibaryons
@@
where the sum is over all the mesons. Note that $\mu=0$ for mesons since the baryon charge is zero for them.
The total (anti)baryon energy density is
$\varepsilon_{B\{\bar{B}\}}=$ $\sum_{i\in B\{\bar{B}\}}\frac{g_{i}}{2\pi^{2}}\int_{0}^{\infty}\>\frac{k^{2}_{i}dk_{i}\tilde{\omega}_{i}}{e^{\big{(}\frac{\omega_{i}-\mu^{i}_{\text{eff}}\{\bar{\mu}_{\text{eff}}\}}{T}\big{)}}+1}~{}$
@@
where $\phi(n)$ represents the energy density correction to avoid double counting the potential. One can determine it by using the condition that $\tilde{\omega}_{i}=\frac{\partial\varepsilon}{\partial n_{i}}$. After doing the derivative of baryon energy density with respect to baryon net number density and using Eq. (26), we get
@@
$P_{B\{\bar{B}\}}(T,\mu)=T\sum_{i\in B\{\bar{B}\}}\frac{g_{i}}{2\pi^{2}}\int_{0}^{\infty}kd^{2}_{i}k_{i}\text{ln}\bigg{[}1+$
@@
To study the thermoelectric coefficients of hot and dense hadron gas in the absence of a magnetic field, we consider the Boltzmann transport equation (BTE) under relaxation time approximation (RTA). Under RTA, the Boltzmann equation can be interpreted as a linear expansion of the single-particle total distribution function ($f_{i}$) around the single-particle equilibrium distribution function ($f_{i}^{0}$). Here, ($f_{i}$) can be written as $f_{i}=f^{0}_{i}+\delta f_{i}$, where ($\delta f_{i}$) represents the deviation from equilibrium. The total single-particle distribution function for $i^{\rm th}$ species at equilibrium, which is given by
@@
where $\omega_{i}=\sqrt{\vec{k_{i}}^{2}+m_{i}^{2}}$ is the single particle energy, $\mu_{B}$ is the baryon chemical potential, $b_{i}$ denotes the baryon number of $i^{\rm th}$ particle, e.g. for baryons $b_{i}=1$, for antibaryons $b_{i}=-1$ and for mesons $b_{i}=0$. The $\pm$ sign stands for fermions and bosons, respectively. In local thermodynamic equilibrium, the spatial dependence of the distribution functions appears due to spatial gradients of temperature and baryon chemical potential. The linearized BTE under RTA in the local rest frame (LRF), for particle species $i$ can be written as Das:2020beh ; Singh:2023pwf ,
@@
where $h=\frac{\varepsilon+P}{n}$ is the enthalpy per particle, $\varepsilon$, $P$, and $n$ are total energy density, total pressure, and net baryon density of the system, respectively. With leading order contribution, we can consider an ansatz of $\delta f_{i}$ as Singh:2023pwf
@@
In the above equation, we have used $\langle v^{l}_{i}v^{j}_{i}\rangle=\frac{1}{3}v_{i}^{2}\delta^{lj}$. Here, the sum is over all the baryons, anti-baryons, and mesons. Also, for a relativistic system, the thermal current is defined with reference to the conserved baryon current. The thermal current arises when energy flows relative to the baryonic enthalpy. Hence, the heat current of the hot hadronic medium can be defined as Gavin:1985ph ,
@@
It is to be noted that mesons contribute through the total enthalpy of the system as well as in the total electrical conductivity of the system, which enters the denominator of the Eq.(51). The presence of a magnetic field makes this picture more complicated, which we discuss in the next section.
@@
To solve the RBTE as given in Eq.(60) we take an ansatz to express the deviation of the distribution function from the equilibrium in the following way,
@@
with $\vec{\Omega}$ being related to a temperature gradient, electric field, the magnetic field and in general, can be written as,
@@
Using Eq.(61) and (62) RBTE as given in Eq.(60) can be expressed as,
@@
$(\Omega_{c_{i}}\tau^{i}_{R})\vec{v}_{i}.(\vec{\nabla T\times B})+(\Omega_{c_{i}}\tau^{i}_{R})^{2}(\vec{\nabla}T.\vec{B})(\vec{v}_{i}.\vec{B})\bigg{]}(-\frac{\partial f^{0}_{i}}{\partial\omega_{i}}).$ (73)

### v2

@@ Abstract
The hot and dense hadronic medium formed during the heavy-ion collisions at the Relativistic Heavy Ion Collider and Large Hadron Collider energies can show thermoelectric effects in the presence of temperature gradients and nonzero baryon chemical potential. In this article, we study the thermoelectric coefficients of an interacting hot and dense hadron gas using the relativistic Boltzmann transport equation under the relaxation time approximation. We discuss the thermoelectric properties within different frameworks of hardon resonance gas models. In the presence of an external magnetic field, the thermoelectric coefficients become anisotropic, which leads to Hall-like thermoelectric coefficients, namely Nernst coefficients, along with the magneto-Seebeck coefficients. For the first time, we also estimate the Thomson coefficient of the medium, which comes into the picture due to the temperature dependence of the Seebeck coefficient of the medium. In the context of studying the thermoelectric generator performance, we calculate the values of the thermoelectric figure of merit of the medium.
@@ I Introduction
In heavy-ion collision experiments such as the Large Hadron Collider (LHC) or the Relativistic Heavy Ion Collider (RHIC), a deconfined state of nuclear matter can be produced at extremely high temperatures ($T$) and/or baryon chemical potentials ($\mu_{B}$). This thermalized state of matter is called quark-gluon plasma (QGP). QGP evolves Brewer:2019oha ; Busza:2018rrf and cools down rapidly, producing a hot and dense hadronic matter that later freezes out leading to the detection of long-lived final-state particles. Studying the hadronic phase followed by the hadronization process is crucial for understanding the full evolution of heavy-ion collisions and extracting hidden information about the properties of dense nuclear matter. Due to the extreme energy density in the initial phase, the system continues to expand and cool down rapidly. Hence, the temperature gradients develop in both space and time. The temperature decreases with increased spatial distance from the collision point, which leads to radial temperature gradients. This temperature gradient plays a significant role in heavy-ion collisions because it, in turn, affects the evolution of the system. Currently, researchers are interested in understanding how the thermoelectric properties of the hot hadron gas influence the evolution of collision dynamics.
@@
The thermoelectric properties refer to the response of a material to a temperature gradient, which can generate an electric potential difference subjected to this temperature gradient book . A baryon current is produced in the medium at a finite baryon chemical potential. Similarly, a charge current will also be produced due to the presence of charges. Due to the baryon and charge currents in the medium, thermal and electrical transport properties are of significant interest in the study of heavy-ion collisions Singh:2023pwf ; Singh:2023ues . The thermoelectric effects are proportional to the derivative of conductivity, and they are generally considered more sensitive to the responses to applied external fields. When thermoelectric coefficients are nonzero, temperature and chemical potential gradients in a conductor can drive an electric current. The Seebeck coefficient measures a material’s efficiency in converting these gradients into electric current and is defined under “open circuit” conditions, where the net current is zero Singh:2024emy . This coefficient has been widely studied in condensed matter systems, including superconductors, quantum dot-coupled Majorana bound states, and various junctions and materials PhysRevB.81.155457 ; PhysRevB.84.075420 ; PhysRevB.101.045101 ; PhysRevB.90.165115 . These thermoelectric transport properties of hot and dense hadron gas have been studied theoretically using various approaches such as relativistic kinetic theory Bhatt:2018ncr ; Das:2020beh ; Das:2021qii , hydrodynamics Greif:2017byw ; Prakash:1993bt ; Kadam:2014xka , and lattice QCD simulations Meyer_2011 ; Kaczmarek_2022 . For some situations, the Seebeck coefficient is not constant but shows a gradient with the temperature gradient. When a current drives through this gradient, the Thomson coefficient can be observed, which describes how much heat is absorbed or released. The amount of heat exchange depends on the medium’s properties and the direction of both the current and the temperature gradient.
@@
In relativistic systems, the thermoelectric formalism differs notably from the nonrelativistic cases: here, thermal current requires a conserved number current, like the baryon number in QCD. This distinction affects the behavior of thermoelectric coefficients, as shown in studies of hadronic matter via the hadron resonance gas (HRG) model Das:2021qii . Studies in isotropic Abhishek:2020wjm and anisotropic Kurian:2021zyb ; Zhang:2020efz QGP systems have also explored this, although using perturbative QCD, which is more applicable at very high temperatures. Near the QCD transition temperature ($T_{c}$), nonperturbative effects are relevant, as the QCD vacuum structure is nontrivial, influenced by chiral symmetry breaking and confinement-related condensates. However, understanding the thermoelectric properties of the hadronic medium is also of equal importance. To understand the hot and dense hadronic matter formed in relativistic heavy-ion collisions, the ideal hadron resonance gas (IHRG) model is a widely used model. It explains the lattice QCD (lQCD) estimations of thermodynamic variables up to $T\approx 150~{}\mathrm{MeV}$ Bellwied:2013cta ; Bellwied:2017ttj , after which the hadrons start to melt, and the disagreements between the IHRG and lQCD appear. Unlike the lQCD, which breaks down at very high $\mu_{B}$ Borsanyi:2013bia ; HotQCD:2014kol , the IHRG model can be used to study highly dense hadronic or nuclear matter. However, the ideal HRG model is unable to explain the higher-order fluctuations of conserved charges, which have been predicted from the lQCD Bazavov:2013dta ; Bazavov:2017dus ; Borsanyi:2018grb . Thus, later on, better alternatives to the simplistic IHRG model have been explored extensively. One such widely used model is the excluded volume hadron resonance gas (EVHRG) model Andronic:2012ut . In this model, the hadrons have some finite radius and thus exert an outward pressure, which is the result of the excluded volume of the hadrons. The EVHRG model explains the thermodynamic properties and the conserved charged fluctuations of the hadronic matter very well, giving a good agreement with the lQCD data Andronic:2012ut ; Kadam:2015xsa ; Kadam:2017iaz ; Pal:2020ucy . There are also other models, such as the repulsive mean-field hadron resonance gas (RMFHRG) model Kapusta:1982qd ; Olive:1980dy , which uses a potential term to introduce the repulsive interaction between the hadrons. Recently, a new model has been explored called the van der Waals hadron resonance gas (VDWHRG) model Vovchenko:2017cbu ; Vovchenko:2016rkn , which takes into account both attractive and repulsive interactions between the hadrons. It has been found that the VDWHRG model is the most reliable HRG model, which agrees with the lQCD estimations Vovchenko:2016rkn ; Samanta:2017yhh . It also gives a second-order liquid-gas phase transition critical point at very high $\mu_{B}$ Vovchenko:2015vxa , which can help study dense nuclear matter present in neutron stars.
@@
In this article, we offer a detailed examination of the thermoelectric properties and associated transport coefficients of the hot and dense hadronic medium in both the presence and the absence of an external magnetic field. We have used four different formalisms, namely, the IHRG, EVHRG, RMFHRG, and VDWHRG, to estimate the thermoelectric coefficients. Further, we also discuss the thermoelectric generator performance of the medium in terms of the thermoelectric figure of merit. This article is organized in the following manner. Section (II) briefly gives the derivation of thermoelectric coefficients with and without an external magnetic field. Section (III) discusses the results in detail and finally, in Sec. (IV), we summarize the study with possible outlooks.
@@ II Formalism
In this section, we discuss the hadron resonance gas model and its extended versions that we use to calculate the thermoelectric coefficients of hot and dense hadron gas. We calculate these coefficients in both the absence and presence of an external magnetic field by using the Boltzmann transport equation under relaxation time approximation. For all our calculations, we have taken all the hadrons and their resonances up to the mass cutoff of 2.6 GeV from the Particle Data Group ParticleDataGroup:2008zun .
@@ II.1 Hadron resonance gas model
II.1.1 Ideal hadron resonance gas model
The IHRG model assumes that a system of hadrons and resonances behaves as an ideal gas, meaning that the particles do not interact with each other and are considered to be point particles. The grand canonical partition function for an ideal hadron resonance gas can be written as Pradhan:2022gbm ,
@@ II.1.1 Ideal hadron resonance gas model
Here, $g_{i}$ is the degeneracy, $k_{i}$ is the momentum, $m_{i}$ is the mass and $\omega_{i}=\sqrt{k_{i}^{2}+m_{i}^{2}}$ is the energy of the $i{\rm th}$ hadron species. The $\pm$ sign corresponds to fermions and bosons, respectively. $\mu_{i}$ denotes the corresponding chemical potential, which is given by
@@
where the baryon chemical potential is given by $\mu_{B}$ and $b_{i}$ denotes the baryon number of the $i{\rm th}$ hadron. Here, we have considered a simple case by taking the vanishing charge and strangeness chemical potential, i.e., $\mu_{Q}$ = $\mu_{S}$ = 0. The pressure $P_{i}$, energy density $\varepsilon_{i}$, and number density $n_{i}$ can now be obtained from the partition function, given as,
@@
II.1.2 Excluded volume hadron resonance gas model
The EVHRG model is an extension of the IHRG model that takes into account the finite size of hadrons. Meanwhile, in an IHRG model, hadrons are treated as pointlike particles without considering their physical volume. The EVHRG model incorporates repulsive interactions by introducing an excluded volume for hadrons, preventing overlap, and thereby providing a more realistic description of hot and dense hadronic matter created in heavy-ion collisions. In thermodynamically consistent excluded volume formulation, the transcendental equation for the pressure can be written as Rischke:1991ke ; Bhattacharyya:2013oya
@@ II.1.2 Excluded volume hadron resonance gas model
II.1.3 van der Waals hadron resonance gas model
@@
The total pressure in the VDWHRG model comes from the mesons, baryons, and antibaryons separately because of the interaction and is given by Samanta:2017yhh ; Vovchenko:2016rkn
@@
Here, $M$, $B$, and $\bar{B}$ represent mesons, baryons, and antibaryons. $\mu^{*M}$ is the modified chemical potential of mesons because of the excluded volume correction ($b_{M}$), and $\mu^{*B}$ and $\mu^{*\bar{B}}$ are the modified chemical potentials of baryons and antibaryons due to VDW interactions ($a$ and $b$Sarkar:2018mbk . The modified chemical potential for mesons and (anti)baryons can be obtained from Eq. (2) and Eq. (15) as,
@@
where $n_{M}$, $n_{B}$ and $n_{\bar{B}}$ are the modified number densities of mesons, baryons and antibaryons, respectively, which are given by
@@
II.1.4 Repulsive mean-field hadron resonance gas model
The RMF corrections are often implemented within a mean-field approximation, where the repulsive interactions are treated as an effective mean field acting on each hadron. This mean field modifies the single-particle energies of the hadrons, affecting their thermodynamic properties such as the equation of state, pressure, and density. A repulsive mean-field approach is utilized to incorporate the effects of repulsive interactions among hadrons, as it was used in Refs. Kapusta:1982qd ; Olive:1980dy and recently, in the case of baryons in Ref.Huovinen:2017ogf . In this approach, it is assumed that the repulsive interactions lead to a shift in the single-particle energy and is given by
$\tilde{\omega}_{i}=\sqrt{\vec{k_{i}}^{2}+m_{i}^{2}}+U(n)=\omega_{i}+U(n),$ (25)
@@ II.1.4 Repulsive mean-field hadron resonance gas model
In this work, we assign different repulsive interaction parameters for baryons and mesons. We denote the mean-field parameter for baryons ($B$) and antibaryons ($\bar{B}$) by $K_{B}$, while for mesons we denote it by $K_{M}$. Thus, for baryons (antibaryons)
@@
where $n_{i}$ is the number density of $i{\rm th}$ hadronic species. Note that $n_{B}$, $n_{\bar{B}}$, and $n_{M}$ are total baryon, antibaryon and meson number densities, respectively. Explicitly, for baryons,
@@
where the sum is over all the baryons. Here, $\mu_{\text{eff}}^{i}=b_{i}\mu_{B}-U(n_{B})$ is the baryon effective chemical potential, with $b_{i}$ being the baryonic number of $i{\rm th}$ baryon and $\mu_{B}$, the baryon chemical potential. Similarly, for antibaryons
@@
where the sum is over all the mesons. Note that $\mu=0$ for mesons since the baryon charge is zero for them. The total (anti)baryon energy density is
$\varepsilon_{B\{\bar{B}\}}=$ $\sum_{i\in B\{\bar{B}\}}\frac{g_{i}}{2\pi^{2}}\int_{0}^{\infty}\>\frac{k^{2}_{i}dk_{i}\tilde{\omega}_{i}}{e^{\big{(}\frac{\omega_{i}-\mu^{i}_{\text{eff}}\{\bar{\mu}^{i}_{\text{eff}}\}}{T}\big{)}}+1}~{}$
$+\phi_{B\{\bar{B}\}}(n_{B\{\bar{B}\}}),$
$=$ $\sum_{i\in B\{\bar{B}\}}\frac{g_{i}}{2\pi^{2}}\int_{0}^{\infty}\>\frac{k^{2}_{i}dk_{i}\tilde{\omega}_{i}}{e^{\big{(}\frac{\tilde{\omega}_{i}-b_{i}\mu^{i}_{B}\{\bar{\mu}^{i}_{B}\}}{T}\big{)}}+1}~{}$
@@
where $\phi(n)$ represents the energy density correction to avoid double-counting the potential. One can determine it by using the condition that $\tilde{\omega}_{i}=\frac{\partial\varepsilon}{\partial n_{i}}$. After doing the derivative of baryon energy density with respect to baryon net number density and using Eq. (26), we get
@@
$P_{B\{\bar{B}\}}(T,\mu)=T\sum_{i\in B\{\bar{B}\}}\frac{g_{i}}{2\pi^{2}}\int_{0}^{\infty}k_{i}^{2}dk_{i}\text{ln}\bigg{[}1+$
@@
To study the thermoelectric coefficients of hot and dense hadron gas in the absence of a magnetic field, we consider the Boltzmann transport equation (BTE) under relaxation time approximation (RTA). Under RTA, the Boltzmann equation can be interpreted as a linear expansion of the single-particle total distribution function ($f_{i}$) around the single-particle equilibrium distribution function ($f_{i}^{0}$). Here, ($f_{i}$) can be written as $f_{i}=f^{0}_{i}+\delta f_{i}$, where ($\delta f_{i}$) represents the deviation from equilibrium. The total single-particle distribution function for $i{\rm th}$ species at equilibrium, is given by
@@
where $\omega_{i}=\sqrt{\vec{k_{i}}^{2}+m_{i}^{2}}$ is the single particle energy, $\mu_{B}$ is the baryon chemical potential, $b_{i}$ denotes the baryon number of $i{\rm th}$ particle, e.g. for baryons $b_{i}=1$, for antibaryons $b_{i}=-1$ and for mesons $b_{i}=0$. The $\pm$ sign stands for fermions and bosons, respectively. In local thermodynamic equilibrium, the spatial dependence of the distribution functions appears due to spatial gradients of temperature and baryon chemical potential. The linearized BTE under RTA in the local rest frame (LRF), for particle species $i$ can be written as Das:2020beh ; Singh:2023pwf ,
@@
where $h=\frac{\varepsilon+P}{n}$ is the enthalpy per particle, $\varepsilon$, $P$, and $n$ are total energy density, total pressure, and net baryon density of the system, respectively. With leading-order contribution, we can consider an ansatz of $\delta f_{i}$ as Singh:2023pwf ; Gavin:1985ph
@@
In the above equation, we have used $\langle v^{l}_{i}v^{j}_{i}\rangle=\frac{1}{3}v_{i}^{2}\delta^{lj}$. Here, the sum is over all the baryons, antibaryons, and mesons. Also, for a relativistic system, the thermal current is defined with reference to the conserved baryon current. The thermal current arises when energy flows relative to the baryonic enthalpy. Hence, the heat current of the hot hadronic medium can be defined as Gavin:1985ph ,
@@
It is to be noted that mesons contribute through the total enthalpy of the system as well as in the total electrical conductivity of the system, which enters the denominator of Eq.  eqref equnew11. The presence of a magnetic field makes this picture more complicated, which we discuss in the next section.
@@
To solve the RBTE as given in Eq.(60) we take an ansatz to express the deviation of the distribution function from the equilibrium in the following way Gavin:1985ph
@@
with $\vec{\Omega}$ being related to a temperature gradient, electric field, the magnetic field and in general, can be written as
@@
Using Eqs.(61) and (62) RBTE as given in Eq.(60) can be expressed as,
@@
$(\Omega_{c_{i}}\tau^{i}_{R})\vec{v}_{i}.(\vec{\nabla T\times B})+(\Omega_{c_{i}}\tau^{i}_{R})^{2}(\vec{\nabla}T.\vec{B})(\vec{v}_{i}.\vec{B})\bigg{\}}\bigg{]}(-\frac{\partial f^{0}_{i}}{\partial\omega_{i}}).$ (73)